It takes Amy twice as long to deliver papers as it does Nancy. How long would it take each girl to deliver the papers by herself if they can deliver the papers together in 32 minutes?

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

It takes Amy twice as long to deliver papers as it does Nancy. How long would it take each girl to deliver the papers by herself if they can deliver the papers together in 32 minutes?
---------------------
Nancy DATA:
time= x min/job ; rate = 1/x job/min
---
Amy DATA:
time = 2x min/job ; rate = 1/2x job/min
----
Together DATA:
time = 32 min/job ; rate = 1/32 job/min
===========================================
Equation:
rate + rate = together rate
1/x + 1/2x = 1/32
Multiply thru by 32x to get:
32 + 16 = x
x = 48 min (time for Nancy to do the job)
---
2x = 96 min (time for Amy to do the job)
===========================================
Cheers,
Stan H.

Answer by richwmiller(17219)

Just curious.
Have you looked for similar problems before posting? Have you looked for lessons before posting?